OpenStudy (anonymous):

Solve for x; 9.2=0.6667log(x) Can anyone solve this equation step by step?

OpenStudy (anonymous):

@Hero @ganeshie8 can you help please?

OpenStudy (nincompoop):

use a calculator

OpenStudy (anonymous):

the calculator I have doesn't provide me with variables. I've tried multiple online calculators and they're not helping me.

OpenStudy (nincompoop):

www.wolframalpha.com

OpenStudy (nincompoop):

x = ab when solving for a, what do you normally do?

hero (hero):

1. Change the decimals to a fraction. (this will ensure a more exact answer in the end) \(9.2 = 9 + \dfrac{2}{10}\) \(0.6667 = \dfrac{2}{3}\) So you have: \(9 + \dfrac{2}{10} = \dfrac{2}{3} \log(x)\)

hero (hero):

\(9 + \dfrac{2}{10} = \dfrac{92}{10}\) so \(\dfrac{92}{10} = \dfrac{2}{3}\log(x)\)

hero (hero):

Next multiply both sides by \(\dfrac{3}{2}\): \(\left(\dfrac{3}{2}\right) \left(\dfrac{92}{10}\right) = \log(x)\)

hero (hero):

The last step is the tricky part. The hint is that \(10^{log(x)} = x\)

OpenStudy (anonymous):

alright, i got it. thanks